# The 230 GHz Variability of Numerical Models of Sagittarius A* II. The Physical Origins of the Variability

### v1

@@ 4 Discussion
Figure 14: The averaged correlation coefficient computed between the total gas internal energy $U_{\rm tot}(r)$ integrated within a volume of radius $r$ and all the $R_{\rm Low}=1$ light curves. The maximum correlation coefficient exceeds $0.6$ for all spins, indicating a moderately positive correlation.
Figure 15: The rescaled time series of $U_{\rm tot}$, and $L_{\gamma}$ that is computed at $R_{\rm Low}=1$. Here, the $U_{\rm tot}$ time series is shown as a black line, corresponding to the point of maximum correlation coefficient shown in Figure 14. $L_{\gamma}$ are all shown as blue dotted lines. The ‘wiggling’ of $U_{\rm tot}$ moderately follow that of $L_{\gamma}$.
In Chan et al. (2024), we conjectured that the high $M_{\Delta T}$ of numerical models is caused by the high sensitivity of the electron temperature prescription (Equation 1) to a slightly changing $\beta$ within $\beta=10^{-1}-10^{1}$. Thus, if one can close the gap between $R_{\rm Low}$ and $R_{\rm High}$, such a sensitivity could be reduced and thus potentially help reduce $M_{\Delta T}$. However, in this study, even though we find that increasing $R_{\rm Low}$ could reduce $M_{\Delta T}$ for black holes with $a>0$, it is merely an optical depth effect — gas becomes optically thick and blocks the photon rings from illuminating. Nonetheless, we do not rule out the possibility that the shape of Equation 1 does play an important role in causing the high $M_{\Delta T}$ when the photon rings are optically thin. Without changing $R_{\rm Low}$, one could easily tune Equation 1 so that the transition between $R_{\rm Low}$ and $R_{\rm High}$ has a flatter slope, and we leave this study for the future.
We also find that increasing $R_{\rm Low}$ would come with a cost of making MAD flux eruptions more visible and help contribute to a higher $M_{\Delta T}$, especially for black holes with $a<0$. However, given that MAD models with $a>0$ are more favoured (Event Horizon Telescope Collaboration et al., 2022), we believe that varying $R_{\rm Low}$ would at least be an intermediate solution to the variability crisis. In the longer term, one should address the origin of the variabilities of the photon rings. We conjecture it might be due to some abnormal amount of heating close to the black hole. Since heating affects the plasma internal energy, to test our idea, we first show the domain-wise modulation index of the gas internal energy $u_{g}=P/(\gamma-1)$ in Figure 13 (a) and (b) We find that the gas internal energy is highly variable close to the black hole and also along the jet wall. The modulation index is also reduced as $a\rightarrow 0$. Next, we compute the total internal energy within a radius $r$ as
@@
Then, given a radius $r$, we compute $U_{\rm tot}$ across the snapshot to form a time series. We then compute the average correlation coefficient between the $U_{\rm tot}$ time series and the $R_{\rm Low}=1$ light curves. This could give us the correlation coefficient versus the radius $r$. In calculating the correlation coefficient, we rescale $U_{\rm tot}$ and the light curves (denote it by $L_{\gamma}$ for convenient discussion) such that $U_{\rm tot}\rightarrow U_{\rm tot}/\langle U_{\rm tot}\rangle$ where bracket denotes the mean value. We do this also for $L_{\gamma}$. This rescaling gives the transformed quantity a standard deviation equal to the modulation index. Hence, the ‘wiggling’ of the transformed quantity reflects its variability. We first show this result in Figure 14. Despite that the correlation coefficients vary differently against $r$ for different black hole spin $a$, the maximum coefficient is $\geq 0.6$, indicating a moderately positive correlation between the $u_{g}$ and the $230$ GHz light curve variability.
Next, we extract the time series of $U_{\rm tot}$ that corresponds to the maximum value with respect to $r$ in Figure 14 and show them in Figure 14. In the same figure, we also append the $230$ GHZ light curves at $R_{\rm Low}=1$. We find that the (rescaled) $U_{\rm tot}$ moderately follow the variation of the (rescaled) $L_{\gamma}$. Note that the radius we considered to compute $U_{\rm tot}$ is within a few $r_{g}$ from the black hole. Thus, it is likely that an abnormal amount of heating near the black hole contributes to the variability of the photon ring at $R_{\rm Low}=1$. The source of heating could be grid-scale dissipation. Nonetheless, it suggests that we should carefully tune the simulation parameters to control the frequency of or released by these heating if one wants to reduce $M_{\Delta T}$. In short, the solution to the variability crisis should contain two ingredients: i) simulations that better control the amount of dissipative heating, and ii) a better electron temperature prescription that controls the response of electron temperature to the heating.
@@ 5 Conclusion
The Event Horizon Telescope Collaboration has successfully observed the galactic centre black hole Sgr A*. Model comparison with observational data provided valuable constraints on the black hole properties, such as the spin, observer inclination, plasma temperature, accretion mode, etc. However, current GRMHD models are inconsistent with the $3$-hour, $230$ GHz flux variability ($M_{\Delta T}$) constraint (Event Horizon Telescope Collaboration et al., 2022), thus creating the variability crisis for modelling Sgr A*. In Chan et al. (2024), we address this problem by varying one of the fixed parameters in the ion-electron temperature ratio prescription functions (Equation 1) — $R_{\rm Low}$, representing the ion-electron temperature ratio at the strongly magnetized regime. We find increasing $R_{\rm Low}$ beyond $1$ helps reduce $M_{\Delta T}$ for almost all black holes with $a>0$, but this is not the case for almost all black holes with $a<0$. The variation of $M_{\Delta T}$ are insensitive to the changes in $R_{\rm Low}$ for black holes with $a=0$. In this study, we aim to address the parameter dependence of $M_{\Delta T}$ on increasing $R_{\rm Low}$ through analyzing suites of GRMHD simulation and GRRT snapshots. We select representative models for each black hole spin and examine the origin of variabilities as seen in the $230$ GHz flux light curves.
@@
First, for black holes with $a>0$, we find that the variability for the $R_{\rm Low}=1$ models originates from the photon rings, which are consistent with our findings in Chan et al. (2024). As $R_{\rm Low}$ increases, the fluid becomes more optically thick, blocking and preventing the photon ring from illuminating, thus reducing the $M_{\Delta T}$. Therefore, reducing $M_{\Delta T}$ due to increasing $R_{\rm Low}$ is merely an optical depth effect. As $R_{\rm Low}$ is further increased, MAD flux eruptions become more visible and help offset the reduced variability of the obstructed photon rings. An optimal solution for the lowest $M_{\Delta T}$ occurs when the variability due to both photon rings and MAD flux eruption events are combined to a minimum. Additionally, we identified and illustrated MHD flux tubes created in MAD flux eruptions in the image domain for the first time.
@@
Second, we observe the same phenomenon for black holes with spin $a<0$. However, the optically thick fluid contains voids that cannot cover the photon rings. Additionally, flux eruptions have a larger spatial scale and are thus more visible. Hence, the increase in the optical depth accompanied by the rise of $R_{\rm Low}$ do not reduce variabilities originating from the photon rings substantially and, on the other hand, enhances variabilities due to MAD flux eruptions. We note that models with $a=-0.94$ and $\theta=179.9$ Deg show a decrease in $M_{\Delta T}$ when one increases $R_{\rm Low}$. We argue that this is a particular case. Since the increasing trend $M_{\Delta T}$ is quite generic across black holes with $a<0$, this particular case does not affect our qualitative understanding of the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$.
@@
Third, the optical depth effect still applies to black holes with spin $a=0$, but their light curves are similar for models with $R_{\rm Low}\neq 1$, and this reveals why $M_{\Delta T}$ is not very sensitive to the variations in $R_{\rm Low}$. We suspect this would be because either we select an interval of snapshots where the black hole is at its quiescent state or that the black hole with $a=0$ is intrinsically quiet. Parameter studies with a longer time interval of the snapshot are required to understand such insensitivity. But given that the current constrains on the spin of Sgr A* prefer $a\neq 0$, we believe that this would not affect our qualitative understanding of the physics of changing $M_{\Delta T}$ on increasing $R_{\rm Low}$, at least within the parameter space of interest.
Fourth, we also briefly discuss the physical origin of the variability of the photon ring at $R_{\rm Low}=1$, the optically thin limit. We suspect an abnormal amount of heating close to the black hole is responsible for producing highly variable emissions. To test our conjecture, we compute the average correlation coefficient between the volume-integrated gas internal energy $U_{\rm tot}$ and the $R_{\rm Low}=1$ light curves. The maximum correlation coefficient is $\geq 0.6$, indicating a moderately positive correlation. The rescaled $U_{\rm tot}$ also moderately follow the variation of the rescaled $L_{\gamma}$. The source of this abnormal amount of heating could be grid-scale dissipation.
Last but not least, we stress that increasing $R_{\rm Low}$ is an intermediate solution that allows us to understand the origin of variabilities through the image domain only. In the longer term, one should address why the photon rings are variable (i.e., identify the source of abnormal heating) and control the amount of energy released, as well as the frequency of the heating. One should also place better constraints on the ion-electron temperature prescriptions and/or electron heating functions to control the response of electron temperature to these heating.
@@
Finally, the current study only considered the effects of $R_{\rm Low}$ as manifested in the intensity domain. Since the reduction in $M_{\Delta T}$ is an optical depth effect, we expect it to be seen through polarimetric signatures. For instance, increasing the optical depth would promote Faraday rotations, increasing the circular and reducing the linear polarization. Comparing resolved or unresolved polarimetric signatures with observations could help test the $R_{\rm Low}\neq 1$ model. Also, a higher optical depth would make the spectral index more positive. Computing either the unresolved spectral index or resolved spectral index map and comparing them with observation could help us constrain the parameter space of $R_{\rm Low}$. We plan to perform these extended studies in our third series of papers.

### v2

@@ 4 Discussion
In Chan et al. (2024), we conjectured that the high $M_{\Delta T}$ of numerical models is caused by the high sensitivity of the electron temperature prescription (Equation 1) to a slightly changing $\beta$ within $\beta=10^{-1}-10^{1}$. Thus, if one can close the gap between $R_{\rm Low}$ and $R_{\rm High}$, such sensitivity could be reduced and potentially help reduce $M_{\Delta T}$. However, in this study, even though we find that increasing $R_{\rm Low}$ could reduce $M_{\Delta T}$ for black holes with $a>0$, it is merely an optical depth effect — gas becomes optically thick and blocks the photon rings from illuminating. Nonetheless, we do not rule out the possibility that the shape of Equation 1 plays an important role in causing the high $M_{\Delta T}$ when the photon rings are optically thin. Without changing $R_{\rm Low}$, one could easily tune Equation 1 so that the transition between $R_{\rm Low}$ and $R_{\rm High}$ has a flatter slope, and we leave this study for the future.
We have shown that the reduction of $M_{\Delta T}$ due to the increase in $R_{\rm Low}$ is more prominent for models with $a>0$, with a few particular cases. It is interesting to examine if such a reduction is systematic across samples, provided that one does not increase $R_{\rm Low}$ excessively. To achieve this, we extract light curves $L_{\gamma}$ for $R_{\rm Low}=1$ and $10$ across all $R_{\rm High}$ and $\theta$, focusing on $a=+0.94$ and $0.5$. Specifically, we perform the transformation $L_{\gamma}\rightarrow L_{\gamma}/\langle L_{\gamma}\rangle$ so that the ‘wiggling’ of the light curve with respect to $1$ represents its time-variability. We present the results in Figure 14. We observed an overall flattening of the sample- mean and range of the transformed light curves, indicating that the effect of reducing $M_{\Delta T}$ when one increases $R_{\rm Low}$ is not a coincidence, at least for models with $a>0$.
@@
Suppose the time variability of the gas internal energy is the major contribution to the time variability of the light curves. In that case, its source is likely plasma heating, which directly impacts the plasma’s internal energy. The source of heating could be grid-scale dissipation. To address the variability crisis with more accurate modeling of plasma heating, one should carefully tune the simulation parameters to control the frequency of, or the energy released by, these heating events, if one aims to reduce $M_{\Delta T}$. The solution should contain two key components: (i) simulations that better control the amount of dissipative heating, and (ii) a better electron temperature prescription that regulates the response of electron temperature to the heating. To better control the amount/frequency of dissipative heating, one could consider using less diffusive numerical schemes or more first-principle modeling of dissipation, such as viscous GRMHD or resistive GRMHD. Additionally, one should look for an electron heating function that does not respond rapidly to changes in $u_{g}$.
Another reason for the high time variability of numerical models could be due to the $\sigma$ cutoff. We show that the source of the light curve time-variability is mostly along the jet funnel wall, while all the emission inside the funnel wall is set to zero by the default $\sigma=1$ cut. If the $\sigma=1$ boundary is highly variable, it is evident that the emission is also highly variable due to the steep step function of emission at $\sigma=1$. If this is the major cause of the variability, one could consider smoothing the emission transition at $\sigma=1$ using an analytic function such as the logistic function. Lastly we should remark that the $M_{3}$ maps we shown in Figure 15 might subject to spatial-variability, and thus might not reflect the time-variability of the lightcurves, which are summation of the all emission and thus should be less sensitive to the spatially varying emission point (i.e., same emission strength but located at different position at different time). Lastly, we note that in GRMHD models, the evolution of the gas internal energy near the $\sigma=1$ interface remains uncertain due to significant density gradients, which could also contribute to the increased variability.
@@ 5 Conclusion
The Event Horizon Telescope Collaboration has successfully observed the galactic center black hole Sgr A*. Model comparison with observational data provided valuable constraints on the black hole properties, such as spin, observer inclination, plasma temperature, accretion mode, etc. However, current GRMHD models are inconsistent with the 3-hour, $230$ GHz flux variability ($M_{\Delta T}$) constraint (Event Horizon Telescope Collaboration et al., 2022), thus creating the variability crisis for modeling Sgr A*. In Chan et al. (2024), we addressed this problem by varying one of the fixed parameters in the ion-electron temperature ratio prescription functions (Equation 1) — $R_{\rm Low}$, representing the ion-electron temperature ratio in the strongly magnetized regime. We found that increasing $R_{\rm Low}$ beyond 1 helps reduce $M_{\Delta T}$ for almost all black holes with $a>0$, but this is not the case for almost all black holes with $a<0$. The variation of $M_{\Delta T}$ is insensitive to changes in $R_{\rm Low}$ for black holes with $a=0$. In this study, we aim to address the parameter dependence of $M_{\Delta T}$ on increasing $R_{\rm Low}$ through analyzing suites of GRMHD simulation and GRRT snapshots. We select representative models for each black hole spin and examine the origin of variabilities as seen in the $230$ GHz flux light curves. Our major findings are:
  1. 1.
@@
    For black holes with $a>0$, we find that the variability for the $R_{\rm Low}=1$ models originates from the photon rings, which is consistent with our findings in Chan et al. (2024). As $R_{\rm Low}$ increases, the fluid becomes more optically thick, blocking and preventing the photon ring from illuminating, thus reducing $M_{\Delta T}$. Therefore, reducing $M_{\Delta T}$ due to increasing $R_{\rm Low}$ is merely an optical depth effect. As $R_{\rm Low}$ is further increased, MAD flux eruptions become more visible and help offset the reduced variability of the obstructed photon rings. An optimal solution for the lowest $M_{\Delta T}$ occurs when the variability due to both photon rings and MAD flux eruption events are combined to a minimum.
  2. 2.
@@
    We observe the same phenomenon for black holes with spin $a<0$. However, the optically thick fluid contains voids that cannot cover the photon rings. Additionally, flux eruptions have a larger spatial scale and are thus more visible. Hence, the increase in the optical depth accompanied by the rise of $R_{\rm Low}$ does not substantially reduce variabilities originating from the photon rings and, on the other hand, enhances variabilities due to MAD flux eruptions. We note that models with $a=-0.94$ and $\theta=179.9$ Deg show a decrease in $M_{\Delta T}$ when $R_{\rm Low}$ is increased. We argue that this is a particular case. Since the increasing trend in $M_{\Delta T}$ is quite generic across black holes with $a<0$, this particular case does not affect our qualitative understanding of the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$.
  3. 3.
@@
    The optical depth effect still applies to black holes with spin $a=0$, but their light curves are similar across models with $R_{\rm Low}\neq 1$, which reveals why $M_{\Delta T}$ is not very sensitive to the variations in $R_{\rm Low}$. We suspect this is because either we selected an interval of snapshots where the black hole is in its quiescent state, or the black hole with $a=0$ is intrinsically quiet. Parameter studies with a longer time interval of snapshots are required to understand such insensitivity. However, given that current constraints on the spin of Sgr A* prefer $a\neq 0$, we believe that this would not affect our qualitative understanding of the physics of changing $M_{\Delta T}$ by increasing $R_{\rm Low}$, at least within the parameter space of interest.
  4. 4.
    We also briefly discuss the physical origin of the variability of the lightcurves at $R_{\rm Low}=1$, the optically thin limit. We show that the origin of emission is predominantly along the jet funnel wall and is also variable at the same position, which is where the gas internal energy is also the most variable. Thus, we suspect that the time variability of gas internal energy is related to that of the light curves. Grid-scale dissipative heating could be the source of the variability of $u_{g}$. We also note that the abrupt $\sigma=1$ cutoff in emission, along with the uncertainty in the evolution of $u_{g}$ within GRMHD models at the $\sigma=1$ interface, could serve as additional sources of variability.
@@
Last but not least, we stress that increasing $R_{\rm Low}$ is an intermediate solution that allows us to understand the origin of variabilities through the image domain only. In the longer term, one should address why the $R_{\rm Low}=1$ models are variable. Finally, the current study only considered the effects of $R_{\rm Low}$ as manifested in the intensity domain. Since the reduction in $M_{\Delta T}$ is an optical depth effect, we expect it to be seen through polarimetric signatures. For instance, increasing the optical depth would promote Faraday rotations, increasing the circular and reducing the linear polarization. Comparing resolved or unresolved polarimetric signatures with observations could help test the $R_{\rm Low}\neq 1$ model. Also, a higher optical depth would make the spectral index more positive. Computing either the unresolved spectral index or resolved spectral index map and comparing them with observation could help us constrain the parameter space of $R_{\rm Low}$. We plan to perform these extended studies in our third series of papers.